if P(A)= 7/36, P(B)= 1/6, and P(A and B)= 0, what can you say about P(A or B)?
Answer by stanbon(75887)

if P(A)= 7/36, P(B)= 1/6, and P(A and B)= 0, what can you say about P(A or B)?
---
P(A or B) = P(A) + P(B) - P(A and B)
P(A or B) = (7/36) + (1/6) - 0
P(A or B) = 7/36 + 6/36 = 13/36
-----------------------------------
Note: Since P(A and B) = 0, A and B are disjoint.
and P(A or B) is simply P(A) + P(B)
=======================================
Cheers,
Stan H.